# Accuracy of the Yee FDTD Scheme for Normal Incidence of Plane Waves on Dielectric and Magnetic Interfaces

### v1

@@ Abstract
This paper analyzes the accuracy of the standard Yee finite-difference time-domain (FDTD) scheme for simulating normal incidence of harmonic plane waves on planar interfaces between lossless, linear, homogeneous, isotropic media. We consider two common FDTD interface models based on different staggered-grid placements of material parameters. For each, we derive discrete analogs of the Fresnel reflection and transmission coefficients by formulating effective boundary conditions that emerge from the Yee update equations. A key insight is that the staggered grid implicitly spreads the material discontinuity over a transition layer of one spatial step, leading to systematic deviations from exact theory. We quantify these errors via a transition-layer model and provide (i) qualitative criteria predicting the direction and nature of deviations, and (ii) rigorous error estimates for both weak and strong impedance contrasts. Finally, we examine the role of the Courant number in modulating these errors, revealing conditions under which numerical dispersion and interface discretization jointly influence accuracy.
@@ 1 Introduction
The finite-difference time-domain (FDTD) method, introduced by Yee [34] and later extended by Taflove & Brodwin [28], is a widely used numerical technique for simulating electromagnetic wave propagation by directly solving Maxwell’s equations in the time domain. It is particularly well-suited for analyzing wave interactions with material interfaces, including attenuation [27], absorption [23], reflection and transmission phenomena [5, 19]. Accurate modeling of electromagnetic wave interactions at material interfaces is crucial in the design of antennas, waveguides, and photonic devices [13]. The FDTD method has become a standard tool [22] for such applications due to its ability to capture transient and steady-state wave behavior in complex media.
@@ 2 Short review of problem
The problem of analyzing the correctness of numerical schemes for PDEs with singularities is not new. One of the earliest works in this area was by Ito & Tanimoto [11]. This work marks one of the earliest practical numerical attempts to handle combined wave diffraction and refraction in coastal zones. Using a finite-difference discretization of a precursor to the mild-slope equation, the authors modeled wave fields around breakwaters over variable bathymetry. Its significance lies in demonstrating that numerical methods could overcome the limitations of purely analytical or graphical techniques in real-world harbor and coastal planning.
@@
Interface-fitted methods are powerful and accurate but geometrically rigid. Their limitations in handling dynamic or complex interfaces motivated the development of immersed and regularization-based strategies—highlighting the ongoing tension in computational science between fidelity, flexibility, and efficiency. Thus, [30] proposed smoothing discontinuous coefficients (e.g., wave speed) via moment-preserving regularization, enabling standard high-order schemes to retain accuracy. This bridges engineering pragmatism, like Ito and Tanimoto’s approach, with modern analysis—offering a lightweight alternative to interface-fitted or immersed methods. The result is a compromise that requires knowledge of interface location but avoids complex stencil modifications.
@@
The present work is devoted to the solution of all the above mentioned problems, as well as to the analysis and generalization of the results obtained. The obtained results are compared for the interfaces of both generalized dielectrics and generalized magnets (moreover, for two extreme cases: both low-contrast and high-contrast interfaces). Additionally, in our work, a more detailed analysis of the accuracy of the obtained expressions was performed. In this regard, both the influence of wavelength discretization and the choice of the Courant number are analyzed.
@@ 3.2 Governing Equations
$\varepsilon_{0}\varepsilon_{\mathrm{r}}\frac{\partial E_{z}}{\partial t}=\frac{\partial H_{y}}{\partial x},\quad\mu_{0}\mu_{\mathrm{r}}\frac{\partial H_{y}}{\partial t}=\frac{\partial E_{z}}{\partial x}.$ (1)
@@
$E_{z}(x,t)=E_{+}\,e^{i(\omega t-kx)}+E_{-}\,e^{i(\omega t+kx)},\quad H_{y}(x,t)=H_{+}\,e^{i(\omega t-kx)}+H_{-}\,e^{i(\omega t+kx)},$ (3)
@@
$k=\omega\frac{n_{\mathrm{r}}}{c}.$ (4)
@@ 3.3 Reflection and Transmission at the Interface
    $E_{z}^{(i)}(x,t)=E_{0}\,e^{i(\omega t-k_{1}x)},\quad H_{y}^{(i)}(x,t)=-\frac{E_{0}}{\eta_{1}}\,e^{i(\omega t-k_{1}x)}.$ (7)
@@
    $E_{z}^{(r)}(x,t)=rE_{0}\,e^{i(\omega t+k_{1}x)},\quad H_{y}^{(r)}(x,t)=r\frac{E_{0}}{\eta_{1}}\,e^{i(\omega t+k_{1}x)}.$ (8)
@@
    $E_{z}^{(t)}(x,t)=tE_{0}\,e^{i(\omega t-k_{2}x)},\quad H_{y}^{(t)}(x,t)=-t\frac{E_{0}}{\eta_{2}}\,e^{i(\omega t-k_{2}x)}.$ (9)
@@ 3.4 Boundary Conditions and Exact Solution
$E_{z}^{(i)}(0,t)+E_{z}^{(r)}(0,t)=E_{z}^{(t)}(0,t),\quad H_{y}^{(i)}(0,t)+H_{y}^{(r)}(0,t)=H_{y}^{(t)}(0,t).$ (11)
@@ 4 Space-time discretization and finite-differences approach
$\frac{\partial\psi(x,t)}{\partial x}\Big|_{(x_{0},t_{0})}=\frac{\psi(x_{0}+\Delta_{x},t_{0})-\psi(x_{0},t_{0})}{\Delta_{x}}+o(1),$ (14)
$\frac{\partial\psi(x,t)}{\partial t}\Big|_{(x_{0},t_{0})}=\frac{\psi(x_{0},t_{0}+\Delta_{t})-\psi(x_{0},t_{0})}{\Delta_{t}}+o(1).$ (15)
@@ 4.1 Yee Algorithm
Let us now refine the approximation (14), (15) using central differences at the spacetime point with coordinates $(x_{0},t_{0})$ shown in Fig. 2 at position $P_{1}$ with discrete indices $m+1/2$ and $q$. The Taylor series expansion of the functions $\psi$ at the points $(x_{0}\pm\Delta_{x}/2,t_{0})$ and $(x_{0},t_{0}\pm\Delta_{t}/2)$ with the accuracy up to the second order of smallness allows to write down
$\frac{\partial\psi(x,t)}{\partial x}\Big|_{(x_{0},t_{0})}=\frac{\psi(x_{0}+\Delta_{x}/2,t_{0})-\psi(x_{0}-\Delta_{x}/2,t_{0})}{\Delta_{x}}+o(h),$ (17)
$\frac{\partial\psi(x,t)}{\partial t}\Big|_{(x_{0},t_{0})}=\frac{\psi(x_{0},t_{0}+\Delta_{t}/2)-\psi(x_{0},t_{0}-\Delta_{t}/2)}{\Delta_{t}}+o(h).$ (18)
Using approximations (17), (18) for point $P_{1}$ yields the FDTD-analogue of the second equation (1) in the following form
@@
By carrying out a similar approximation procedure (17), (18) for the time-space grid node $P_{2}$, from the first equation (1) one can obtain the FDTD-analogue of the Ampère-Maxwell law equation
@@
The equations (19) and (21) obtained above represent a coupled system of equations that allows the implementation of a recurrent scheme for determining all components of the electromagnetic field using known initial conditions. For this reason, that in the literature on the FDTD method these equations are known as the Update Equations [26, 10].
@@
is optimal for modeling the electromagnetic properties of a homogeneous, non-dispersive medium. Here, $n_{\mathrm{r}}$ is the relative refractive index of the medium
$n_{\mathrm{r}}\equiv\sqrt{\varepsilon_{\mathrm{r}}\mu_{\mathrm{r}}}.$ (23)
@@ 4.2 Generation of the Incident Wave Field
From the equations (19), (21) it is obvious that in the framework of the Condition 2 accepted by us, nontrivial dynamics of the electromagnetic field is possible only in two cases [21]:
@@
Here $f^{q}[m]$—is the finite-difference analog of the continuous electric component $E_{z}^{(i)}(x,t)$ of the incident electromagnetic field (7).
@@ 4.3 FDTD Analogs of Incident, Reflected and Transmitted Wave Fields
For future reference, we also list here the FDTD analogs of the fields (7) – (9). Here and everywhere else in the following, we assume that the interface between Medium 1 and Medium 2 (plane $x=0$) is the Yee grid node with the number $m=b$ (all necessary details will be set out below in § 5).
@@
We note separately that the magnitude $E_{0}$ and cyclic frequency $\omega$ of a harmonic incident wave is assumed to be a given quantity for which there is no difference between the continuous and discretized versions. At the same time, the FDTD-impedance matches with its continuous counterpart, the proof of which can be found in [26, 7.5]. Because of that, this quantities writting in (26) – (28) and below without tilde sign.
@@ 4.4 FDTD Dispersion
The last thing we need to discuss before moving directly to achieving the goal of this study is the consideration of numerical dispersion, which inevitably takes place in FDTD simulations (even in the case of non-dispersive and homogeneous materials). However, since this issue is studied in detail in the article [20] (also, see [26, 7.4]), here we will reproduce only the main result, which we will need later.
@@ 5.1 Necessary definitions and lemmas
The equality (36) follows directly from the first equation of the system (1) after applying the approximation (17), (18) for the node of the Yee grid with coordinates $m$, $q+1/2$, labeled in Fig. 2 as point $P_{2}$, with the using of Definitions 1 and 2.
@@
The same actions performed for the second equation (1) at the point $P_{1}$ of Fig. 2 lead to the equation (37). ∎
@@ 5.2 Dielectrics
Consider the interface of two dielectrics with relative permittivities $\varepsilon_{1}$ and $\varepsilon_{2}$. Let the relative permeability of both media be the same and equal to $\mu$ (for generality of the problem formulation; $\mu$ is not necessarily equal to $1$). Geometry of this model is shown in Figure 3. On that scheme we suppose, that Medium 1 has an extend of $b$ nodes, such that
@@ 5.3 Magnetics
Similarly (40), this expression was obtained after substituting the fields (26) – (28) into the corresponding equation (11) and canceling $E_{0}e^{i\omega q\Delta_{t}}$.
@@
Substituting now (51) – (53) into (50), reducing the common multiplier $E_{0}e^{i\omega q\Delta_{t}}$, multiplying by $\eta_{1}\Delta_{x}$ and grouping the terms with coefficient $\widetilde{t}$ in the right-hand side of the equality, we finally obtain the equation (48). ∎
@@ 6.1 Dielectric/dielectric interface
The real part of the denominator of (60) The real part of the denominator of eqThrm3-Step6 equation 1 e-ik1x2 2 eik2x2 1 k1x2 2 k2x2 equation coincides
@@
coincides with the denominator of the fraction, defining $\widetilde{t}$ in (54). Therefore, it remains only to prove that the imaginary part of this denominator is zero. This follows from the dispersion relation (29) in Lemma 1.
@@
is identically zero. This finally confirms that the equation (60) is the first equality (54). To complete the proof, it remains to verify that substituting $\widetilde{t}$ from (54) into (40) does indeed lead to the second equality (54). ∎

### v2

@@ Abstract
This paper analyzes the accuracy of the standard Yee finite-difference time-domain (FDTD) scheme for simulating normal incidence of harmonic plane waves on planar interfaces between lossless, linear, homogeneous, isotropic media. We consider two common FDTD interface models based on different staggered-grid placements of material parameters. For each, we derive discrete analogs of the Fresnel reflection and transmission coefficients by formulating effective boundary conditions that emerge from the Yee update equations. A key insight is that the staggered grid implicitly spreads the material discontinuity over a transition layer of one spatial step, leading to systematic deviations from exact theory. We quantify these errors via a transition-layer model and provide (i) qualitative criteria predicting the direction and nature of deviations, and (ii) rigorous error estimates for both weak and strong impedance contrasts. Finally, we examine the role of the Courant number in modulating these errors, revealing conditions under which numerical dispersion and interface discretization jointly influence accuracy. This analysis provides quantitative error estimates that are directly applicable to simulation practice, offers a transition-layer interpretation that bridges classical FDTD with modern immersed-interface methods, and establishes benchmarks for validating structure-preserving discretizations of Maxwell’s equations.
@@ 1 Introduction
The finite-difference time-domain (FDTD) method, introduced by Yee [38] and later extended by Taflove and Brodwin [32], is a widely used numerical technique for simulating electromagnetic wave propagation by directly solving Maxwell’s equations in the time domain. It is particularly well-suited for analyzing wave interactions with material interfaces, including attenuation [30], absorption [27], reflection and transmission phenomena [5, 24]. Accurate modeling of electromagnetic wave interactions at material interfaces is crucial in the design of antennas, waveguides, and photonic devices [16]. The FDTD method has become a standard tool [26] for such applications due to its ability to capture transient and steady-state wave behavior in complex media.
@@
1.1 Relevance of Yee FDTD analysis in the modern computational landscape
While significant advances have been made in the last two decades in the numerical solution of Maxwell’s equations—notably through finite-element exterior calculus (FEEC) [1, 2] and geometric numerical integration methods [25, 10]—the Yee FDTD scheme remains a cornerstone of computational electromagnetics for several compelling reasons.
  1. 1.
    Industrial and educational ubiquity. The Yee scheme is implemented in virtually every commercial and open-source electromagnetic solver (e.g., Lumerical, CST, MEEP [26]), and it is the first time-domain method taught in most electromagnetics curricula worldwide. Rigorous error analysis of this foundational method therefore has immediate practical impact: it enables practitioners to interpret simulation results with quantified confidence and to design discretizations that meet application-specific accuracy requirements.
  2. 2.
    Benchmarking and validation. Modern structure-preserving methods (FEEC, geometric integrators) are often validated against the Yee scheme in canonical configurations. Our explicit error estimates for Fresnel coefficients at material interfaces provide a quantitative benchmark for such comparisons, clarifying which deviations arise from interface discretization versus bulk numerical dispersion.
  3. 3.
    Conceptual insight: the transition-layer interpretation. A key contribution of this work is the explicit identification of the staggered-grid interface as an implicit transition layer of width $\Delta_{x}$ (Fig. 6). This perspective bridges the gap between the classical Yee discretization and modern immersed-interface or regularization-based approaches [19, 34], suggesting pathways for hybrid methods that retain Yee’s simplicity while improving interface accuracy.
  4. 4.
    Foundational role for multiscale and multilayer systems. The one-dimensional normal-incidence configuration analyzed here is not merely a simplification: it is the building block for transfer-matrix methods, multilayer coating design, and homogenization of periodic structures. Our error estimates therefore propagate directly to these applied contexts, where rapid parameter sweeps often rely on 1D FDTD kernels for efficiency.
  5. 5.
    Pedagogical value for structure-preserving discretizations. By deriving discrete Fresnel coefficients directly from the Yee update equations, we illustrate how conservation properties (e.g., energy balance $R+T=1$) emerge—or fail to emerge—from a staggered-grid discretization. This concrete example can serve as an accessible entry point for students and researchers transitioning from classical FDTD to more abstract structure-preserving frameworks.
In summary, while FEEC and geometric integration represent powerful theoretical advances, the Yee scheme’s practical dominance, pedagogical centrality, and role as a validation baseline ensure that rigorous accuracy analysis of its interface treatment remains both timely and impactful. This work contributes precisely such an analysis, with explicit error estimates that are immediately applicable to simulation practice and informative for the development of next-generation methods.
@@ 2 Short review of problem
The problem of analyzing the correctness of numerical schemes for PDEs with singularities is not new. One of the earliest works in this area was by Ito and Tanimoto [14]. This work marks one of the earliest practical numerical attempts to handle combined wave diffraction and refraction in coastal zones. Using a finite-difference discretization of a precursor to the mild-slope equation, the authors modeled wave fields around breakwaters over variable bathymetry. Its significance lies in demonstrating that numerical methods could overcome the limitations of purely analytical or graphical techniques in real-world harbor and coastal planning.
@@
Interface-fitted methods are powerful and accurate but geometrically rigid. Their limitations in handling dynamic or complex interfaces motivated the development of immersed and regularization-based strategies—highlighting the ongoing tension in computational science between fidelity, flexibility, and efficiency. Thus, Tornberg and Engquist [34] proposed smoothing discontinuous coefficients (e.g., wave speed) via moment-preserving regularization, enabling standard high-order schemes to retain accuracy. This bridges engineering pragmatism, like Ito and Tanimoto’s approach, with modern analysis—offering a lightweight alternative to interface-fitted or immersed methods. The result is a compromise that requires knowledge of interface location but avoids complex stencil modifications.
@@
The present work is devoted to the solution of all the above mentioned problems, as well as to the analysis and generalization of the results obtained. The obtained results are compared for the interfaces of both generalized dielectrics and generalized magnets (moreover, for two extreme cases: both low-contrast and high-contrast interfaces). Additionally, in our work, a more detailed analysis of the accuracy of the obtained expressions was performed. In this regard, both the influence of wavelength discretization and the choice of the Courant number are analyzed. By explicitly connecting our error estimates to practical simulation parameters (wavelength discretization $N_{\lambda}$, Courant number $S_{\mathrm{c}}$) and to the broader context of structure-preserving numerical methods, we aim to make this analysis both mathematically rigorous and immediately useful to practitioners.
@@ 3.2 Governing Equations
$\varepsilon_{0}\varepsilon_{\mathrm{r}}\frac{\partial E_{z}}{\partial\mathrm{t}}=\frac{\partial H_{y}}{\partial x},\quad\mu_{0}\mu_{\mathrm{r}}\frac{\partial H_{y}}{\partial\mathrm{t}}=\frac{\partial E_{z}}{\partial x}.$ (1)
@@
$E_{z}(x,\mathrm{t})=E_{+}\,e^{i(\omega\mathrm{t}-kx)}+E_{-}\,e^{i(\omega\mathrm{t}+kx)},\quad H_{y}(x,\mathrm{t})=H_{+}\,e^{i(\omega\mathrm{t}-kx)}+H_{-}\,e^{i(\omega\mathrm{t}+kx)},$ (3)
@@
$k=\omega\frac{n_{\mathrm{r}}}{c},$ (4)
where $n_{\mathrm{r}}$ is the relative refractive index of the medium
$n_{\mathrm{r}}\equiv\sqrt{\varepsilon_{\mathrm{r}}\mu_{\mathrm{r}}}.$ (5)
@@ 3.3 Reflection and Transmission at the Interface
    $E_{z}^{(i)}(x,\mathrm{t})=E_{0}\,e^{i(\omega\mathrm{t}-k_{1}x)},\quad H_{y}^{(i)}(x,\mathrm{t})=-\frac{E_{0}}{\eta_{1}}\,e^{i(\omega\mathrm{t}-k_{1}x)}.$ (8)
@@
    $E_{z}^{(r)}(x,\mathrm{t})=rE_{0}\,e^{i(\omega\mathrm{t}+k_{1}x)},\quad H_{y}^{(r)}(x,\mathrm{t})=r\frac{E_{0}}{\eta_{1}}\,e^{i(\omega\mathrm{t}+k_{1}x)}.$ (9)
@@
    $E_{z}^{(t)}(x,\mathrm{t})=tE_{0}\,e^{i(\omega\mathrm{t}-k_{2}x)},\quad H_{y}^{(t)}(x,\mathrm{t})=-t\frac{E_{0}}{\eta_{2}}\,e^{i(\omega\mathrm{t}-k_{2}x)}.$ (10)
@@ 3.4 Boundary Conditions and Exact Solution
$E_{z}^{(i)}(0,\mathrm{t})+E_{z}^{(r)}(0,\mathrm{t})=E_{z}^{(t)}(0,\mathrm{t}),\quad H_{y}^{(i)}(0,\mathrm{t})+H_{y}^{(r)}(0,\mathrm{t})=H_{y}^{(t)}(0,\mathrm{t}).$ (12)
@@ 4 Space-time discretization and finite-differences approach
$\frac{\partial\psi(x,\mathrm{t})}{\partial x}\Big|_{(x_{0},\mathrm{t}_{0})}=\frac{\psi(x_{0}+\Delta_{x},\mathrm{t}_{0})-\psi(x_{0},\mathrm{t}_{0})}{\Delta_{x}}+o(1),$ (15)
$\frac{\partial\psi(x,\mathrm{t})}{\partial\mathrm{t}}\Big|_{(x_{0},\mathrm{t}_{0})}=\frac{\psi(x_{0},\mathrm{t}_{0}+\Delta_{t})-\psi(x_{0},\mathrm{t}_{0})}{\Delta_{t}}+o(1).$ (16)
@@ 4.1 Yee Algorithm
Let us now refine the approximation (15), (16) using central differences at the spacetime point with coordinates $(x_{0},\mathrm{t}_{0})$ shown in Fig. 2 at position $P_{1}$ with discrete indices $m+1/2$ and $q$. The Taylor series expansion of the functions $\psi$ at the points $(x_{0}\pm\Delta_{x}/2,\mathrm{t}_{0})$ and $(x_{0},\mathrm{t}_{0}\pm\Delta_{t}/2)$ with the accuracy up to the second order of smallness allows to write down
$\frac{\partial\psi(x,\mathrm{t})}{\partial x}\Big|_{(x_{0},\mathrm{t}_{0})}=\frac{\psi(x_{0}+\Delta_{x}/2,\mathrm{t}_{0})-\psi(x_{0}-\Delta_{x}/2,\mathrm{t}_{0})}{\Delta_{x}}+o(\Delta_{x}),$ (18)
$\frac{\partial\psi(x,\mathrm{t})}{\partial\mathrm{t}}\Big|_{(x_{0},\mathrm{t}_{0})}=\frac{\psi(x_{0},\mathrm{t}_{0}+\Delta_{t}/2)-\psi(x_{0},\mathrm{t}_{0}-\Delta_{t}/2)}{\Delta_{t}}+o(\Delta_{t}).$ (19)
Using approximations (18), (19) for point $P_{1}$ yields the FDTD-analogue of the second equation in (1) in the following form
@@
By carrying out a similar approximation procedure (18), (19) for the time-space grid node $P_{2}$, from the first equation in (1) one can obtain the FDTD-analogue of the Ampère-Maxwell law equation
@@
The (20) and (22) obtained above represent a coupled system of equations that allows the implementation of a recurrent scheme for determining all components of the electromagnetic field using known initial conditions. For this reason, that in the literature on the FDTD method these equations are known as the Update Equations [31, 13].
@@
is optimal for modeling the electromagnetic properties of a homogeneous, non-dispersive medium.
@@ 4.2 Generation of the Incident Wave Field
From the (20), (22) it is obvious that in the framework of the Condition 2 accepted by us, nontrivial dynamics of the electromagnetic field is possible only in two cases [23]:
@@
Here $f^{q}[m]$—is the finite-difference analog of the continuous electric component $E_{z}^{(i)}(x,\mathrm{t})$ of the incident electromagnetic field (8).
@@ 4.3 FDTD Analogs of Incident, Reflected and Transmitted Wave Fields
For future reference, we also list here the FDTD analogs of the fields (8) – (10). Here and everywhere else in the following, we assume that the interface between Medium 1 and Medium 2 (plane $x=0$) is the Yee grid node with the number $m=b$ (all necessary details will be set out below in section 5).
@@
We note separately that the magnitude $E_{0}$ and cyclic frequency $\omega$ of a harmonic incident wave is assumed to be a given quantity for which there is no difference between the continuous and discretized versions. At the same time, the FDTD-impedance matches with its continuous counterpart, the proof of which can be found in [31, 7.5]. Because of that, these quantities are writting in (26) – (28) and below without tilde sign.
@@ 4.4 FDTD Dispersion
The last thing we need to discuss before moving directly to achieving the goal of this study is the consideration of numerical dispersion, which inevitably takes place in FDTD simulations (even in the case of non-dispersive and homogeneous materials). However, since this issue is studied in detail in the article [22] (also, see [31, 7.4]), here we will reproduce only the main result, which we will need later.
@@ 5.1 Necessary definitions and lemmas
The equality (36) follows directly from the first equation of the system (1) after applying the approximation (18), (19) for the node of the Yee grid with coordinates $m$, $q+1/2$, labeled in Fig. 2 as point $P_{2}$, with the use of Definitions 1 and 2.
@@
The same actions performed for the second equation in (1) at the point $P_{1}$ of Fig. 2 lead to the (37). ∎
@@ 5.2 Dielectrics
Consider the interface of two dielectrics with relative permittivities $\varepsilon_{1}$ and $\varepsilon_{2}$. Let the relative permeability of both media be the same and equal to $\mu$ (for generality of the problem formulation $\mu$ is not necessarily equal to $1$). Geometry of this model is shown in Fig. 3. On that scheme we suppose, that Medium 1 has an extend of $b$ nodes, such that
@@ 5.3 Magnetics
Similarly (40), this expression was obtained after substituting the fields (26) – (28) into the (12) and canceling $E_{0}e^{i\omega q\Delta_{t}}$.
@@
Substituting now (51) – (53) into (50), reducing the common multiplier $E_{0}e^{i\omega q\Delta_{t}}$, multiplying by $\eta_{1}\Delta_{x}$ and grouping the terms with coefficient $\widetilde{t}$ in the right-hand side of the equality, we finally obtain the (48). ∎
@@ 6.1 Dielectric/dielectric interface
The real part of the denominator of (60)
@@
coincides with the denominator of the fraction, defining $\widetilde{t}$ in (54). Therefore, it remains only to prove that the imaginary part of this denominator is zero. This follows from the dispersion relation (29) in Lemma 1
@@
is identically zero. This finally confirms that the (60) is the first equality (54). To complete the proof, it remains to verify that substituting $\widetilde{t}$ from (54) into (40) does indeed lead to the second equality (54). ∎